OpenStudy (icalibear):

Solve -2 + x/12 = 5/6

OpenStudy (icalibear):

does x = 34

OpenStudy (icalibear):

does x = 34

OpenStudy (anonymous):

x=34

OpenStudy (icalibear):

yeah i dont think you make it into a decimal you have to x by 12 to cancel the fraction

OpenStudy (anonymous):



OpenStudy (icalibear):

thanks

OpenStudy (anonymous):

or we could have just found the greatest common factor which is 12 and multiplied everything throughout

OpenStudy (anonymous):

12( -2 + x/12= 5/6) -24+ x = 10 x=34